# Strong Screening Rules for Group-based SLOPE Models

### v1

@@ F.3.2 Additional results for the logistic model
Figure A11: The number of groups/variables in $\mathcal{E},\mathcal{A}$ as a function of the regularization path for the logistic model with SGS and gSLOPE, shown for different values of the correlation ($\rho$) and $p$. The results are averaged over $100$ repetitions, with $95\%$ confidence intervals shown.
Figure A12: Runtime (in seconds) for screening against no screening as a function of $p$, broken down into different correlation cases, for the logistic model. The results are averaged over $100$ repetitions, with standard errors shown.
@@
Method $p$ Runtime (s)
Runtime (s) screen
@@
SGS $500$ $118\pm 1$ $113\pm 1$ $28\pm 1$ $33\pm 1$ $38\pm 1$ $8\pm 0.2$ $6\pm 0.3$ $29\pm 1$ $8\times 10^{-9}\pm 7\times 10^{-10}$
SGS $1625$ $248\pm 2$ $538\pm 9$ $50\pm 2$ $59\pm 2$ $64\pm 2$ $11\pm 0.4$ $7\pm 1$ $63\pm 3$ $5\times 10^{-9}\pm 1\times 10^{-9}$
SGS $2750$ $374\pm 2$ $868\pm 12$ $85\pm 3$ $104\pm 2$ $115\pm 3$ $21\pm 1$ $7\pm 0.4$ $57\pm 2$ $3\times 10^{-9}\pm 5\times 10^{-10}$
SGS $3875$ $558\pm 4$ $1280\pm 19$ $116\pm 3$ $148\pm 3$ $164\pm 3$ $30\pm 1$ $8\pm 0.4$ $54\pm 1$ $2\times 10^{-9}\pm 2\times 10^{-10}$
SGS $5000$ $737\pm 5$ $1498\pm 19$ $141\pm 4$ $188\pm 3$ $209\pm 4$ $40\pm 1$ $8\pm 0.4$ $54\pm 1$ $1\times 10^{-9}\pm 2\times 10^{-10}$
F.4 Additional results from the real data analysis
@@
Table A7: Variable screening metrics for SGS applied to real data in 5.2. The number of variables in $\mathcal{A}_{v},\mathcal{S}_{v},\mathcal{E}_{v}$, and $\mathcal{K}_{v}$ are shown. The results are averaged across the nine pathway collections, with standard errors shown.
Dataset Method Input
$\operatorname{card}(\mathcal{A}_{v})$ $\operatorname{card}(\mathcal{S}_{v})$ $\operatorname{card}(\mathcal{E}_{v})$ $\operatorname{card}(\mathcal{K}_{v})$
SGS Cancer $5651$ $71\pm 1$ $211\pm 4$ $212\pm 4$ $1\pm 0.04$
SGS Colitis $10259$ $25\pm 1$ $208\pm 6$ $209\pm 6$ $1\pm 0.07$
@@
Table A8: General and group screening metrics for SGS and gSLOPE applied to real data in 5.2. General metrics: the runtime (in seconds) for screening against no screening, the number of fitting iterations for screening against no screening (with the number of occasions of failed convergence given in brackets), and the $\ell_{2}$ distance between the fitted values obtained with screening and no screening. Group screening metrics: the number of groups in $\mathcal{A}_{g},\mathcal{S}_{g},\mathcal{E}_{g}$, and $\mathcal{K}_{g}$. The results are averaged across the nine pathway collections, with standard errors shown.
Method Dataset Input
@@
gSLOPE Cancer $592$ $66\pm 1$ $120\pm 2$ $120\pm 2$ $0\pm 0$ $1719\pm 63(16)$ $4453\pm 103(186)$ $7\times 10^{-8}\pm 7\times 10^{-9}$
gSLOPE Colitis $655$ $32\pm 1$ $56\pm 2$ $56\pm 2$ $0\pm 0$ $1215\pm 73(28)$ $3904\pm 127(140)$ $3\times 10^{-7}\pm 5\times 10^{-8}$
SGS Cancer $592$ $50\pm 1$ $167\pm 2$ $115\pm 2$ $67\pm 0.8$ $72\pm 3(0)$ $127\pm 4(0)$ $1\times 10^{-8}\pm 1\times 8^{-10}$
SGS Colitis $655$ $18\pm 0.6$ $136\pm 4$ $87\pm 3$ $61\pm 2$ $156\pm 9(0)$ $1348\pm 62(0)$ $4\times 10^{-8}\pm 6\times 10^{-9}$

### v2

@@ F.3.2 Additional Results for the Logistic Model
Figure A12: The number of groups/variables in $\mathcal{E},\mathcal{A}$ as a function of the regularization path for the logistic model with SGS and gSLOPE, shown for different values of the correlation ($\rho$) and $p$. The results are averaged over $100$ repetitions, with $95\%$ confidence intervals shown. This figure is the equivalent of Figure A10 for the logistic model.
Figure A13: Runtime (in seconds) for screening against no screening as a function of $p$, broken down into different correlation cases, for the logistic model. The results are averaged over $100$ repetitions, with standard errors shown. This figure is the equivalent of Figure 4 for the logistic model.
@@
Method $p$ Runtime
Runtime screen (s)
@@
SGS $500$ $118\pm 1$ $113\pm 1$ $28\pm 1$ $33\pm 1$ $38\pm 1$ $6\pm 0.3$ $29\pm 1$ $8\times 10^{-9}\pm 7\times 10^{-10}$
SGS $1625$ $248\pm 2$ $538\pm 9$ $50\pm 2$ $59\pm 2$ $64\pm 2$ $7\pm 1$ $63\pm 3$ $5\times 10^{-9}\pm 1\times 10^{-9}$
SGS $2750$ $374\pm 2$ $868\pm 12$ $85\pm 3$ $104\pm 2$ $115\pm 3$ $7\pm 0.4$ $57\pm 2$ $3\times 10^{-9}\pm 5\times 10^{-10}$
SGS $3875$ $558\pm 4$ $1280\pm 19$ $116\pm 3$ $148\pm 3$ $164\pm 3$ $8\pm 0.4$ $54\pm 1$ $2\times 10^{-9}\pm 2\times 10^{-10}$
SGS $5000$ $737\pm 5$ $1498\pm 19$ $141\pm 4$ $188\pm 3$ $209\pm 4$ $8\pm 0.4$ $54\pm 1$ $1\times 10^{-9}\pm 2\times 10^{-10}$
F.4 Data Description
  • carbotax: Carbotax study of ovarian tumor growth.
    • Response (continuous): Relative tumor volume ($\log_{2}$ scale).
    • Data matrix: Gene expression measurements. $10000$ factors were randomly sampled from a collection of $34964$.
    • Grouping structure: Variables are grouped using k-means clustering (Lloyd, 1982).
  • scheetz: Gene expression data in the mammalian eye.
    • Response (continuous): Gene expression measurements for the Trim32 gene.
    • Data matrix: Gene expression measurements for other genes.
    • Grouping structure: Variables are grouped using k-means clustering (Lloyd, 1982).
  • adenoma: Transcriptome profile data to identify formation of colorectal adenomas.
    • Response (binary): Binary labels for whether sample came from adenoma or normal mucosa.
    • Data matrix: Transcriptome profile measurements.
    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.111gsea-msigdb.org/gsea/msigdb/human/collections.jsp. Accessed 08/2024.
  • cancer: Breast cancer patients treated with tamoxifen for 5 years.
    • Response (binary): Binary labels classifying whether the cancer had recurred.
    • Data matrix: Gene expression data.
    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.11footnotemark: 1
  • celiac: Gene expression data of primary leucocytes to classify celiac disease.
    • Response (binary): Binary labels as to whether a patient has celiac disease.
    • Data matrix: Gene expression measurements from the primary leucocytes.
    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.11footnotemark: 1
  • colitis: Blood cells data for classifying whether a patient has colitis.
    • Response (binary): Binary labels classifying whether a patient has colitis.
    • Data matrix: Gene expression measurements.
    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.11footnotemark: 1
  • tumour: Gene expression data of pancreative cancer samples to identify tumorous tissue.
    • Response (binary): Binary labels indicating if a sample is from tumour tissue.
    • Data matrix: Gene expression measurements.
    • Grouping structure: Genes are assigned to pathways (groups) using the C3 regulatory target gene sets.11footnotemark: 1
Table A8: Dataset information for the six datasets used in the real data analysis.
Dataset $p$ $n$ $m$ Group sizes Type Source
carbotax $10000$ $101$ $100$ $[1,126]$ Linear Koussounadis et al. (2014)222downloaded from https://iowabiostat.github.io/data-sets/. Accessed 08/2024.
scheetz $18975$ $120$ $379$ $[1,165]$ Linear Scheetz et al. (2006)0
adenoma $17661$ $64$ $1849$ $[1,646]$ Logistic Sabates-Bellver et al. (2007)333downloaded from https://www.ncbi.nlm.nih.gov/. Accessed 09/2024.
cancer $7057$ $60$ $1277$ $[1,292]$ Logistic Ma et al. (2004)0
celiac $14294$ $132$ $1666$ $[1,570]$ Logistic Heap et al. (2009)0
colitis $11999$ $127$ $1528$ $[1,497]$ Logistic Burczynski et al. (2006)0
tumour $17661$ $52$ $1849$ $[1,646]$ Logistic Pei et al. (2009); Ellsworth et al. (2013); Li et al. (2016)0
F.5 Additional Results from the Real Data Analysis
Figure A14: The ratio of no screen time to screen time ($\uparrow$) of SLOPE, gSLOPE, and SGS applied to the real datasets, for fitting $100$ path models, split into response type. The horizontal grey line represents no screening improvement.
Figure A15: The ratio of the fitting set ($\mathcal{E}_{v}$) to the input dimensionality ($p$) ($\downarrow$) of SLOPE, gSLOPE, and SGS applied to the real datasets, for fitting $100$ path models, split into response type.
Figure A16: The ratio of the fitting set ($\mathcal{E}_{v}$) to the active set ($\mathcal{A}_{v}$) ($\downarrow$) of SLOPE, gSLOPE, and SGS applied to the real datasets, for fitting $100$ path models, split into response type.
Figure A17: The proportion of variables in $\mathcal{S}_{v}$ relative to $p$ for group-only and bi-level screening applied to SGS, plotted along the regularization path for the carbotax, scheetz, adenoma, colitis, and tumour datasets.
@@ F.5 Additional Results from the Real Data Analysis
Table A9: Variable screening metrics for SGS applied to real data in Section 6.2. The number of variables in $\mathcal{A}_{v},\mathcal{S}_{v},\mathcal{E}_{v}$, and $\mathcal{K}_{v}$ are shown. The results are averaged across the nine pathway collections, with standard errors shown.
Method Dataset $\operatorname{card}(\mathcal{A}_{v})$ $\operatorname{card}(\mathcal{S}_{v})$ $\operatorname{card}(\mathcal{E}_{v})$ $\operatorname{card}(\mathcal{K}_{v})$
SGS carbotax $144\pm 6$ $361\pm 10$ $364\pm 10$ $2\pm 0.50$
SGS scheetz $612\pm 74$ $1908\pm 174$ $1915\pm 175$ $7\pm 1.28$
SGS adenoma $2174\pm 23$ $2451\pm 4$ $2463\pm 5$ $12\pm 0.22$
SGS cancer $86\pm 3$ $273\pm 7$ $273\pm 7$ $1\pm 0.05$
SGS celiac $175\pm 10$ $545\pm 22$ $546\pm 22$ $1\pm 0.07$
SGS colitis $68\pm 4$ $409\pm 7$ $411\pm 7$ $2\pm 0.18$
SGS tumour $128\pm 5$ $1189\pm 59$ $1200\pm 59$ $12\pm 0.54$
@@
Table A10: General and group screening metrics for SGS and gSLOPE applied to real data in Section 6.2. General metrics: the runtime (in seconds) for screening against no screening, the number of fitting iterations for screening against no screening (with the number of occasions of failed convergence given in brackets), and the $\ell_{2}$ distance between the fitted values obtained with screening and no screening. Group screening metrics: the number of groups in $\mathcal{A}_{g},\mathcal{S}_{g},\mathcal{E}_{g}$, and $\mathcal{K}_{g}$. The results are averaged across the nine pathway collections, with standard errors shown.
Method Dataset Runtime
Runtime screen (s)
@@
gSLOPE carbotax $865$ $3456$ $71\pm 3$ $124\pm 2$ $124\pm 2$ $0.05\pm 0.05$ $2367\pm 176(1)$ $6433\pm 303(10)$ $1\times 10^{-9}\pm 2\times 10^{-10}$
gSLOPE scheetz $4731$ $7735$ $153\pm 6$ $212\pm 6$ $212\pm 6$ $0\pm 0$ $6252\pm 313(20)$ $7183\pm 332(38)$ $8\times 10^{-10}\pm 8\times 10^{-10}$
gSLOPE adenoma $415$ $12610$ $71\pm 6$ $77\pm 4$ $77\pm 4$ $0\pm 0$ $3956\pm 378(14)$ $7864\pm 394(78)$ $9\times 10^{-6}\pm 1\times 10^{-6}$
gSLOPE cancer $109$ $5003$ $46\pm 3$ $81\pm 5$ $81\pm 5$ $0\pm 0$ $764\pm 74(0)$ $6250\pm 404(40)$ $4\times 10^{-7}\pm 8\times 10^{-8}$
gSLOPE celiac $205$ $7530$ $34\pm 75$ $62\pm 9$ $62\pm 9$ $0\pm 0$ $834\pm 177(0)$ $4764\pm 428(27)$ $4\times 10^{-6}\pm 1\times 10^{-5}$
gSLOPE colitis $148$ $11567$ $69\pm 5$ $114\pm 8$ $114\pm 8$ $0\pm 0$ $749\pm 62(0)$ $9090\pm 206(72)$ $4\times 10^{-6}\pm 7\times 10^{-7}$
gSLOPE tumour $222$ $33405$ $24\pm 1$ $83\pm 4$ $83\pm 4$ $0\pm 0$ $892\pm 67(0)$ $9224\pm 218(86)$ $4\times 10^{-7}\pm 5\times 10^{-8}$
SGS carbotax $63$ $2067$ $40\pm 1$ $138\pm 1$ $119\pm 1$ $168\pm 60(0)$ $3572\pm 350(3)$ $8\times 10^{-10}\pm 3\times 10^{-10}$
SGS scheetz $358$ $5476$ $105\pm 4$ $203\pm 5$ $186\pm 6$ $10\pm 4(0)$ $5039\pm 360(28)$ $7\times 10^{-8}\pm 2\times 10^{-8}$
SGS adenoma $1234$ $6143$ $601\pm 6$ $789\pm 1$ $735\pm 1$ $1172\pm 69(0)$ $4224\pm 338(20)$ $6\times 10^{-7}\pm 5\times 10^{-8}$
SGS cancer $151$ $211$ $70\pm 2$ $295\pm 3$ $182\pm 4$ $124\pm 15(0)$ $147\pm 14(0)$ $2\times 10^{-8}\pm 2\times 10^{-9}$
SGS celiac $339$ $2035$ $120\pm 7$ $460\pm 10$ $288\pm 10$ $55\pm 12(0)$ $1259\pm 112(0)$ $9\times 10^{-5}\pm 2\times 10^{-5}$
SGS colitis $271$ $1821$ $52\pm 3$ $368\pm 3$ $230\pm 3$ $54\pm 10(0)$ $1411\pm 188(0)$ $1\times 10^{-6}\pm 3\times 10^{-7}$
SGS tumour $781$ $7451$ $94\pm 3$ $733\pm 10$ $479\pm 14$ $33\pm 7(0)$ $6006\pm 287(23)$ $7\times 10^{-7}\pm 2\times 10^{-7}$